# Privacy-Preserving Dataset Combination

### v2

@@ 4.5.1 Underlying Score Limitations
4.5.2 Sometimes, Not All Underlying Data Is Needed
Relatedly, when seeing a few samples can successfully identify useful candidate hospitals, $\pi_{p}$ does not always outperform $\pi_{s}$ on small samples.
@@ 4.5.2 Sometimes, Not All Underlying Data Is Needed
Figure 3: Hospital 243: Underlying KL-score performs poorly
Figure 4: Hospital 420: Underlying KL-score performs well
Figure 5: Left: $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ outperforms $\pi_{p}(k=30)$ and $\pi_{0}$. Middle: All strategies perform similarly. Right: $\pi_{p}(k=300)$ outperforms $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$. In all panels, the bars represent the standard deviation.
4.5.3 Encrypted Computation Limitations
Engineering a secure system for machine learning requires both machine learning and software engineering knowledge. We share our our code and method, but also note potential challenges with deploying our secure computation:
@@ 4.5.3 Encrypted Computation Limitations
    Operational: engineering personnel limitations. While our implementation requires little cryptographic knowledge to deploy, it still needs technically-trained staff at each participating hospital to collaborate and maintain.
@@
    Engineering: Extending any MPC protocol is non-trivial, as security engineering is a specialized skill. While $\mathrm{SecureKL}$ applies broadly to other underlying scores in multi-party setups, every new algorithm requires software engineering - prototyping, tuning, debugging —which can be especially costly for hospitals.
@@
    Framework Limitation: While CrypTen is designed to accommodate PyTorch, it is a research tool where not all plain text functionalities are implemented. Writing optimizers – such as L-BFGS – and custom operators that are not readily available requires both machine learning and cryptography knowledge.
@@
    Inherent to Secure Computation: When the method requires significant hyper-parameter tuning, such as using SGD on small batch data with learning rate schedules, plaintext tuning may not transfer perfectly. As detailed in Appendix  8, our hyperparameters for SGD are indeed different in encrypted and plaintext settings. However, encrypted computation hides loss curves and training details by default, complicating development.
@@ 5 Related Work
5.1 Data Valuation and Pricing
The question of how to assess the impact and therefore worth of data has been well studied. Data valuation as a field seeks to quantify the contribution of individual data points or datasets to model performance. Shapley value-based approaches provide theoretically grounded valuations but scale poorly to large datasets [45]. More efficient methods include influence functions [46] and leave-one-out testing [47]. Recent work has extended these concepts to dataset-level valuation [48]. As practitioners create larger datasets, the emergence of data marketplaces has sparked interest in data pricing mechanisms [49]. Query-based pricing [50] and outcome-driven valuations [51] aim to balance seller compensation with buyer utility. While these approaches inform fair data exchange, they typically assume direct access to data, unlike our privacy-preserving method.
5.2 Alternative Approaches to Data Sharing
Recent work has explored several approaches to mitigate data sharing constraints while maintaining model performance. We discuss two primary directions: synthetic data generation and transfer learning from public pretraining.
Synthetic data generation
@@ 5.2 Alternative Approaches to Data Sharing
Synthetic data generation has emerged as a promising approach to expand training datasets while preserving privacy. Generative adversarial networks (GANs) have shown success in generating realistic cancer incidence data [52], medical imaging data [53], and electronic health records [54]. These methods can preserve statistical properties of the original data while providing differential privacy guarantees. Transforming data into a similar form that de-sensitizes certain attributes can be desirable [55, 56, 57, 58, 59]. Yet, to still preserve the utility of the dataset transformed for analytics or learning tasks is challenging by itself [60]. Additionally, outside of the scope of sensitive data that is transformed, little privacy guarantee is available, leading to re-identification risks [61, 60].
@@
However, evaluation of synthetic medical data reveals challenges in capturing rare conditions and maintaining consistent relationships between multiple health variables [52]. For tabular data, methods like CTGAN [62] and TVAE [62] have demonstrated ability to learn complex distributions while preserving correlations between features. However, these approaches often struggle with high-dimensional data and can introduce subtle biases that impact downstream model performance [63]. Recent work has also explored combining synthetic data with differential privacy to provide formal privacy guarantees [64]. While these methods offer stronger privacy protection, they often face significant utility loss, particularly for rare but important cases in the original dataset [65].
Public pre-training and private fine-tuning
Transfer learning via public pretraining has become increasingly popular for domains with limited private data access. BioBERT [17] and ClinicalBERT [15] demonstrated that pretraining on PubMed abstracts and clinical notes can improve performance on downstream medical tasks. Similar approaches have emerged in other regulated domains, including FinBERT [66] for financial applications. However, the effectiveness of transfer learning depends heavily on domain alignment. One study showed that continued pretraining on domain-specific data significantly outperforms generic pretraining when domains differ substantially [16]. This presents challenges for highly specialized fields where public data may not capture domain-specific patterns [67]. Recent work has explored methods to quantify and optimize domain adaptation. Adaptive pretraining strategies [68] and domain-specific vocabulary augmentation [16] have shown promise in bridging domain gaps. However, these approaches still require substantial compute and may not fully capture specialized domain knowledge present in private datasets.
5.3 Benefits of Data Scaling Beyond Performance
Recent work has demonstrated that increasing dataset size and diversity yields benefits beyond raw performance metrics. Large-scale training data has been shown to improve model robustness to distribution shifts [23] and reduce demographic performance disparities [69]. Studies of vision models trained on increasingly large datasets show improved out-of-distribution generalization [14]. Similarly, language models trained on diverse data demonstrate better performance across different domains and demographic groups [70].
5.4 Secure and Confidential Computation
Secure and confidential computation encompasses cryptographic techniques that protect information privacy during computation. In a two-party setup between a model owner and data owner, these methods enable computing joint functions on private inputs without revealing them to other parties.
This requires an encoding scheme $\mathrm{Enc}(\cdot)$ that satisfies the homomorphic property: $\mathrm{Enc}(A)\circ\mathrm{Enc}(B)=\mathrm{Enc}(A\circ B)$, where $A$ and $B$ represent data held by two parties. The inverse function $\mathrm{Enc}^{-1}(\cdot)$ must exist to decode the final output: $\mathrm{Enc}^{-1}(A\circ B)=A\circ B$. Considering an “honest-but-curious” threat model, where parties aim to jointly compute on privately-held data, two main approaches emerge.
Fully Homomorphic Encryption (FHE)
FHE enables arbitrary additions and multiplications on encrypted inputs. While it represents the gold standard for encrypted computation, adapting it to modern machine learning is challenging due to computational constraints from growing ciphertext size. FHE implementations typically use lattice-based schemes requiring periodic “bootstrapping” (key refreshing and noise reduction through re-encryption) via methods like the CKKS scheme [71]. This introduces cryptographic parameters that non-experts struggle to configure effectively.
Secure Multi-party Computation (S-MPC)
SMPC enables multiple parties to compute functions over private inputs while revealing only the final output [33, 34]. An MPC system uses key exchanges, encryption schemes, and secure communication to ensure only encrypted data leaves owner control [72]. Though generally faster than FHE, SMPC’s engineering complexity and communication overhead can limit adoption. Traditional private training approaches that completely hide data can also impede essential model development tasks like inspection, monitoring, and debugging.
5.5 Secure Data Combination
@@
Recent work has explored methods for securely combining datasets while preserving privacy and improving model performance. Early approaches focused on using secure multi-party computation to enable multiple parties to jointly train models without sharing raw data [73]. However, these methods often struggled with computational overhead and communication costs when dealing with large-scale datasets [74]. More recent techniques have introduced frameworks for evaluating potential data partnerships before commitment. These approaches use privacy-preserving protocols to estimate the compatibility and complementarity of different datasets [75, 76]. Some methods focus specifically on measuring distribution shifts between datasets without revealing sensitive information [77]. Several systems have been developed to facilitate secure data combination in specific domains. In healthcare, methods have been proposed for securely combining patient records across institutions while maintaining HIPAA compliance [78]. Financial institutions have explored similar approaches for combining transaction data while preserving client confidentiality [66].
5.6 Other Privacy-Preserving Methods
Federated Learning.
@@ 5.6 Other Privacy-Preserving Methods
Cross-silo federated, decentralized, and collaborative machine learning  [74, 79, 80, 81] focus on acquiring more data through improved data governance and efficient system design. Healthcare machine learning is considered especially suitable, as health records are often isolated [82, 83, 84]. Yet, even though no raw data is shared, model parameters or gradients flow through the system. As the federated computing paradigm offer no privacy guarantee, the system is vulnerable to model inversion [85] and gradients leakage attacks [86, 87]. A subtle but urgent concern is that privacy risks discourage the very formation of the federation when optimisation is traded off with privacy [88, 89]. Building on the insight that useful data is often disparately owned, we tackle the specific incentive problem between pairs of data players where one side trains the model, instead of scaling up a federation (number of parties) to address data access issues. We thus focusing on making this exchange efficient, accurate, and private.
@@
Compared to vanilla Federated Learning, an MPC system [34, 33, 90, 35] provides stronger guarantee in terms of input security. Model owners and data owners can potentially federate their proprietary data, including model weights, training, and testing data, can work together under stringent privacy requirements. Our work extends the line of works  [44, 91, 90] that demonstrates the potential of incorporating MPC in various federated scenarios. On the practical side, unlike mobile-based networks for secure federated learning protocols [80], our system assumes a smaller number of participants, where communication cost and runtime are not dominant concerns.
Differential Privacy
Differential privacy (DP) [92] offers formal privacy guarantees for sharing data and training machine learning models. While DP mechanisms can protect individual privacy when releasing model outputs or aggregated statistics, they face significant limitations for inter-organizational data sharing. The primary challenge is that DP operates on already-pooled data, but organizations are often unwilling to share their raw data in the first place [43]. Even when organizations are willing to share data, the privacy guarantees of DP come at a substantial cost to utility, particularly in machine learning applications. DP-SGD, the standard approach for training deep neural networks with differential privacy, significantly degrades model performance compared to non-private training [93]. This performance impact is especially pronounced in data-constrained settings, where recent work has shown that large models rely heavily on memorization of rare examples that DP mechanisms tend to obscure [94]. The privacy-utility trade-off becomes even more challenging when dealing with high-dimensional data or complex learning tasks. Studies have demonstrated that achieving meaningful privacy guarantees while maintaining acceptable model performance requires prohibitively large datasets [95]. This limitation is particularly problematic in specialized domains like healthcare, where data is inherently limited and performance requirements are stringent [96]. Recent work has attempted to improve the privacy-utility trade-off through advanced composition theorems and adaptive privacy budget allocation [97]. However, these approaches still struggle to match the performance of non-private training, especially when working with modern deep learning architectures [98]. While differential privacy offers important theoretical guarantees, our work focuses on the practical challenge of enabling data owners to evaluate potential partnerships before sharing any data, addressing a key barrier to collaboration that DP alone cannot solve.
@@ 6 Conclusion
Our work demonstrates that privacy-preserving data valuation can help organizations identify beneficial data partnerships while maintaining data sovereignty. Through SecureKL, we show that entities can make informed decisions about data sharing without compromising privacy or requiring complete dataset access. As the AI community continues to grapple with data access challenges, particularly in regulated domains like healthcare, methods that balance privacy and utility will become increasingly critical for responsible advancement of the field. As noted in Section 4.5, our approach has several limitations, including the fact that, despite impressive aggregate results, our method is less effective for individual hospitals; this finding is fertile ground for future work. Additionally, our work present opportunities for follow-up research. Our method assumes static datasets and may not generalize well to scenarios where data distributions evolve rapidly over time. A sequential version of our framework may more closely model dynamic data collaborations. Future work should explore extending these techniques to handle more complex data types and dynamic distribution shifts while maintaining strong privacy guarantees.

### v3

@@ 6.3 SecureKL Error Analysis
Sometimes, Not All Underlying Data Is Needed
Relatedly, when seeing a few samples can successfully identify useful candidate hospitals, $\pi_{s}$ (which is on small samples) outperforms $\pi_{p}$ (which is on full samples).
@@
7 Discussion
7.1 SecureKL Contribution
After establishing that $\pi_{p}$ with $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$is a robust strategy in practical downstream performance, we hereby summarize the benefits of $\mathrm{SecureKL}$ and elaborate on their practical implications.
Matching Plaintext Performance in Downstream Tasks. Our major contribution is to match plaintext performance with no data sharing. Using MPC provides input privacy, meaning that if both hospitals only want to know the resulting score, the computation can be done without leaking original data. This strong guarantee can significantly ease the tension related to privacy and compliance in setting up a collaboration, leading to a practical ”data appraisal stage” in data-limited high stakes domains.
Gain from Data Availability. In contrast to limited-sample approaches, a key advantage for our method $\pi_{p}$ is that it takes advantage of all the underlying data – generally impossible with non-secure methods for private data in heavily regulated domains. The general intuition is that data is localized; therefore, once a good target hospital is identified, we would prefer to acquire all of the data. It may be tempting to assert that we prefer the highest $k$ for data addition algorithms as well. In our experiments, while this is generally true, the smaller $k$ sometimes outperform larger $k$ in plaintext strategy $\pi_{s}$, which we investigate in Section 6.3 and in Figure 9. This occasionally non-monotonic behavior mirrors the challenge of data combination itself: even within one source dataset for the same estimator, more data is not necessarily better. This suggests the potential for a hospital-specific alternative to sharing a large amount of data for some source hospital, and points to future directions to using secure computation on a minimal-sized sample dataset for minimal performance overhead while remaining private.
Potential Improvements to SecureKL In the case where that output can be sensitive, i.e., when a source hospital queries a target hospital multiple times and accrues information through the score function, the output can also be made privacy-preserving through differentially private data releases, such as using randomized response (Dwork et al., 2014).
7.2 Potential Challenges to Broader Adoption
Our code is readily usable by small organizations. While our approach generalizes to model-based measures (by substituting $g_{\mathrm{SKL}}$) and scales to multiple parties, our work also uncovered deployment limitations.
@@ 7.2 Potential Challenges to Broader Adoption
    Operational: engineering personnel limitations. While our implementation requires little cryptographic knowledge to deploy, it still needs technically-trained staff at each participating hospital to collaborate and maintain. This skill is similar to using pre-packaged software, cleaning data, and setting up network calls.
@@
    Engineering Extensions: Extending any MPC protocol is non-trivial, as security engineering is a specialized skill. While $\mathrm{SecureKL}$ applies broadly to other underlying scores in multi-party setups, validating a new MPC algorithm requires software engineering – prototyping, tuning, debugging – and numerical verification – akin to data analytics and research - likely requiring technical talents who can be especially costly for hospitals to retain in-house.
@@
    Framework Limitation: While CrypTen is designed to accommodate PyTorch, it is a research tool where not all plain text functionalities are implemented. For example, writing optimizers – such as L-BFGS – and custom operators that are not readily available requires both machine learning and cryptography knowledge. Moreover, the protocol incurs additional computational overhead, especially if hyper-parameters become more complex to sweep [^4] For our work, the performance metrics are provided in Appendix 14 for reference. This will likely improve with time, as new frameworks address design shortcomings.
@@
    Inherent to Secure Computation: When the method requires significant hyper-parameter tuning, such as using SGD on small batch data with learning rate schedules, plaintext tuning may not transfer perfectly. As detailed in Appendix 12, our hyperparameters for SGD differ in encrypted and plaintext settings. Thus, as encrypted computation hides loss curves and training details by default, development is expected to be complex. This is because both hospitals want to ensure model fit with secure evaluation, but may not want to expend the computational cost of private hyperparameter sweeping.
Figure 9: Mean AUC ($\pm$ standard deviation). Left: $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$ outperforms $\pi_{s}(k=30)$ and $\pi_{0}$. Middle: All strategies perform similarly. Right: $\pi_{s}(k=300)$ outperforms $\mathrm{Secure}\mathrm{KL}_{\mathcal{X}\mathcal{Y}}$.
@@ 8 Related Work
We briefly relate alternative approaches towards dataset privacy for sharing. More work on data pricing, differential privacy, and private fine-tuning are included in Appendix 16.2.
Augmenting existing data with synthetic data in medical domains
@@
Synthetic data generation has emerged as a promising approach to expand training datasets while preserving privacy. Generative adversarial networks (GANs) have shown success in generating realistic cancer incidence data (Goncalves et al., 2020), medical imaging data (Thambawita et al., 2022), and electronic health records (Baowaly et al., 2019). These methods preserve statistical properties of the original data while providing differential privacy guarantees. Transforming data into a similar form that desensitizes certain attributes can be desirable (Drechsler, 2011; Howe et al., 2017; Nikolenko, 2021; Gonzales et al., 2023; Sweeney, 2002). Yet, to still preserve the utility of the dataset transformed for analytics or learning tasks is challenging by itself (Jordon et al., 2021). Additionally, outside the scope of sensitive data that is transformed, little privacy guarantee is available, leading to re-identification risks (Narayanan and Shmatikov, 2006; Jordon et al., 2021).
@@
In addition, evaluation of synthetic medical data reveals challenges in capturing rare conditions and maintaining consistent relationships between multiple health variables (Goncalves et al., 2020). For tabular data, methods like CTGAN and TVAE (Xu et al., 2019) have demonstrated ability to learn complex distributions while preserving correlations between features. However, these approaches often struggle with high-dimensional data and can introduce subtle biases that impact downstream model performance (Assefa et al., 2020). Recent work has also explored combining synthetic data with differential privacy to provide formal privacy guarantees (Jordon et al., 2018). While these methods offer stronger privacy protection, they often face significant utility loss, particularly for rare but important cases in the original dataset (Yang et al., 2024a).
Secure Data Combination
@@
Recent work has explored methods for securely combining datasets while preserving privacy and improving model performance. Early approaches focused on using secure multi-party computation to enable multiple parties to jointly train models without sharing raw data (Aono et al., 2017). However, these methods often struggled with computational overhead and communication costs when dealing with large-scale datasets (McMahan et al., 2017). More recent techniques have introduced frameworks for evaluating potential data partnerships before commitment. These approaches use privacy-preserving protocols to estimate the compatibility and complementarity of different datasets (Leung et al., 2019; Chakraborty et al., 2024). Some methods focus specifically on measuring distribution shifts between datasets without revealing sensitive information (Duan et al., 2021). Others trained the downstream model in private, but limited to LASSO (van Egmond et al., 2021). Several systems have been developed to facilitate secure data combination in specific domains. In healthcare, methods have been proposed for securely combining patient records across institutions while maintaining HIPAA compliance (Raisaro et al., 2018; van Egmond et al., 2021). Financial institutions have explored similar approaches for combining transaction data while preserving client confidentiality (Liu et al., 2021).
Federated Learning
@@
Cross-silo federated, decentralized, and collaborative ML (McMahan et al., 2017; Li et al., 2020; Bonawitz et al., 2019; Kairouz et al., 2021) focus on acquiring more data through improved data governance and efficient system design. Healthcare machine learning is considered especially suitable, as health records are often isolated (Rieke et al., 2020; Xu et al., 2021; Nguyen et al., 2022; Cho et al., 2025). Yet, even though no raw data is shared, model parameters or gradients flow through the system. As the federated computing paradigm offers no privacy guarantee, the system is vulnerable to model inversion (Geiping et al., 2020) and gradients leakage attacks (Boenisch et al., 2023; Zhu et al., 2019). A subtle but urgent concern is that privacy risks discourage the very formation of the federation when optimization is traded off with privacy (Lyu et al., 2022; Raynal and Troncoso, 2024). Building on the insight that useful data is often disparately-owned, we tackle the specific incentive problem between pairs of players where one side trains the model, instead of scaling up the number of parties through a federation.
@@
Compared to vanilla Federated Learning, an MPC system (Shamir, 1979; Yao, 1982; Bonawitz et al., 2017; Knott et al., 2021) provides stronger guarantee in terms of input security. Model owners and data owners can potentially federate their proprietary data, including model weights, training, and testing data, can work together under stringent privacy requirements. Our work extends the line of work by  (Xu et al., 2022; Yang et al., 2024b; Bonawitz et al., 2017) that demonstrates the potential of incorporating MPC in various federated scenarios. On the practical side, unlike mobile-based networks for secure federated learning protocols (Bonawitz et al., 2019), our system assumes a smaller number of participants, where communication cost and runtime are not dominant concerns.
Comparing with Federated Learning with Privacy Guarantees
Privacy incentivizes federation (Usynin et al., 2024). Specifically, preserving privacy between the parties under federated learning uses secure computation methods (Truex et al., 2019; Bonawitz et al., 2017; Stripelis et al., 2021; Cho et al., 2025; Froelicher et al., 2021) combined with differential privacy (Xu et al., 2023; Usynin et al., 2024; Avent et al., 2017), trusted execution environments (Mo et al., 2021) – an approach that is coined ”Privacy-in-Depth” in Kairouz et al. (2021).
Performing dataset evaluation with MPC, as in SecureKL, can be seen as an extension of data federation, by adding a separate privacy-preserving component before all parties commit to an entire federated learning system. It is low-commitment, as privacy is preserved by default, and there is no requirement to continue in any system; it tackles the incentives’ problem blocking collaboration. By demonstrating utility, the two parties foster trust. Our work complements the line of ambitious systems that incorporate MPC in potential federated scenarios where the participants are semi-honest (Yang et al., 2024b; Bonawitz et al., 2017; Cho et al., 2025).
@@ 9 Conclusion
Our work demonstrates that privacy-preserving data valuation can help organizations identify beneficial data partnerships while maintaining data sovereignty. Through SecureKL, we show that entities can make informed decisions about data sharing without compromising privacy or requiring complete dataset access. As the AI community continues to grapple with data access challenges, particularly in regulated domains like healthcare, methods that balance privacy and utility will become increasingly critical for responsible advancement of the field. As noted in Section 6.3, our approach has several limitations, including the fact that, despite impressive aggregate results, our method is less effective for individual hospitals, which motivates future work. Our method assumes static datasets and may not generalize well to scenarios where data distributions evolve rapidly over time. A sequential version of our framework may more closely model dynamic data collaborations. Future work should explore extending these techniques to handle more complex data types and dynamic distribution shifts while maintaining strong privacy guarantees.